OpenStudy (anonymous):

how to write x^2=16-z^2 in cylindrical and spherical coordinates??

OpenStudy (experimentx):

for cylindrical coordinates. z = z x = rcos\theta y = rsin\theta http://en.wikipedia.org/wiki/Cylindrical_coordinate_system#Cartesian_coordinates put those values in place of x,y,z then you will have it.

OpenStudy (experimentx):

for spherical coordinates, z = rcos\theta x = r sin\theta * cos\phi y = r sin\theta * sin\phi

OpenStudy (anonymous):

means we can't find the value of r right?just replace x=r cos theta into the equation...and the new equation contains r,theta and z...am i right?

OpenStudy (experimentx):

yes we can find the value of r, r = sqrt(x^2+y^2), and theta = arctan(y/x) but these are going to be our new variables instead of x,y, z

OpenStudy (anonymous):

what type of surface we get from equation (y^2/25)-(x^2/4)=z/3??

OpenStudy (experimentx):

loks like hyperboloid

OpenStudy (anonymous):

how you know the surface is hyperbolic paraboloid based on the equation?

OpenStudy (anonymous):

The surface is a cylinder around the y-axis with radius 4

OpenStudy (experimentx):

google + wolfram plus (y^2/25)-(x^2/4) looks like hyperbola.

OpenStudy (anonymous):

The quation is \[ x^2 + z^2 = 4^2 \] Any plan perpendiula to the y-axis cus th surfacei ircle of radus 4

OpenStudy (anonymous):

Any plan perpendiular to the y-axis cus thr surface in a circle of radus 4

OpenStudy (experimentx):

yes it is, but how do we do for (y^2/25)-(x^2/4) = z/3

OpenStudy (anonymous):

But, we are not talking about this equation now. We are talking about \[ x^2 + z^2=16 \]

OpenStudy (experimentx):

yeah ... I understand that.

OpenStudy (mendicant_bias):

(lol)